OpenStudy (thadds2003):

Medal and fan. Josh has been running for 6 min and completes a lap in 7 min. Gabe has been running for 2 min and completes a lap in 9 min. How many more laps must Josh and Gabe complete to have run the same amount of time?

OpenStudy (thadds2003):

@twister02

OpenStudy (thadds2003):

@ronlover101

OpenStudy (thadds2003):

@readergirl12

OpenStudy (thadds2003):

@Mehek14

OpenStudy (thadds2003):

@Hero

OpenStudy (thadds2003):

@malcolmmcswain

OpenStudy (ronlover101):

Josh's speed is 1 lap/7 min Gabe's speed is 1 lap/9 min Josh has covered 6/7 of a lap Gabe has covered 2/9 of a lapthe answer is 2 or So j completed 6/7 of a lap gabe has completed 2/9 lap so 1 j lap=7 mins s o 7j j=num of laps 1 g lap=9 mins so 9g g=num of laps then 6+j and g+2 so when will the times be equal 7j+6=9g+2 subtract 2 from both sides 7j+4=9g divide both sides by 9 7/9j+4/9=g so to get a whole number of laps, we must subsitute a number when it is multiplied by 7 then add 4, we get a multiplue of 9 so if j=2 7/9(2)+4/9=g=14/9+4/9=g=18/9=g=2 so josh and gabe must run 2 lap each because 6+(2 times 7)=2+(9 times 2) 6+14=2+18 20=20 true so the answer is 2 laps http://brainly.com/question/300623

OpenStudy (thadds2003):

@pizza12

OpenStudy (ronlover101):

Do you get what I was saying @Thadds2003 ?